OpenStudy (anonymous):

the graph of y = x^3 -9x^2 - 16x +1 has a slope of 5 at two points. Find the coordinates of the points

OpenStudy (bahrom7893):

y'=3x^2-18x-16=5

OpenStudy (bahrom7893):

solve for x

OpenStudy (anonymous):

I know that y'= 3x^2-18x-16 0=3x^2-18-21 I am stuck

OpenStudy (turingtest):

you have forgotten how to solve the quadratic\[3x^2-18x-21=0\]this is definately a skill you need to have down for calculus first divide out the common factor of 3, then see if you can factor the equation if not, use the quadratic formula

OpenStudy (anonymous):

(3x-21)(x+1)..Is this correct

OpenStudy (turingtest):

yeah, but you could factor out the 3\[3x^2-18x-21=0\implies x^2-6x-7=0\]\[\implies(x-7)(x+1)=0\]

OpenStudy (anonymous):

x=7 x=-1 and i plugged it in the org equation..Thanks